Mission Plan – November 27th

There are 4 main results we want to achieve by the end of the rotation:
Operate for 12 days according to the MARS-V analog station training program. Individually jot down all comments and improvements.
Make 3 freeze-dried recipes everyday for 2 weeks. Individually jot down all comments and improvements.
Take the MARS-V analog Mars suit and MDRS suit on every EVA. Individually jot down all comments and improvements.
Learn all that we can from MDRS.

EVA Report – December 1st

Sol #5: A Journey of Discovery to Somerville Overlook

EVA Crew Lead: Journalist Sunny
Team Members: HSE Officer Duluu, Geologist Davaa, Engineer Muggi
CapCom: Tunga
Objective: Our mission was to reach and explore Compass Rock, utilizing drone technology for remote reconnaissance and surveying areas inaccessible by foot or rover.

Weather Conditions: The weather was favorable, with cloudy skies and temperatures ranging from 0 to 5 degrees Celsius, conducive to outdoor exploration activities.

Activities and Observations:

The EVA began at 10:15 a.m., with our team navigating the terrain in the Spirit and Opportunity rovers.
We successfully reached Somerville Overlook, which presented a landscape strikingly similar to the Martian surface.
Our drone operations were a highlight, capturing stunning images and providing valuable insights into areas beyond our physical reach.
During the EVA, we observed a faster-than-expected depletion of the rovers’ batteries. This led to a strategic pause at approximately halfway, with the rovers at about 55% battery capacity.
Upon returning to base at 12:45 p.m., we noted the rovers still had 30% battery left, an intriguing finding that may influence future planning.
Challenges and Solutions:

Battery Concerns: The unexpected battery drain posed a potential risk to our mission, but careful management and monitoring allowed for a safe return with sufficient power reserves.
Communication Disruption: We experienced a loss of connection with CapCom Tungaa over a long distance. This was promptly reported to Mission Support, who advised patience until the connection was restored, which occurred upon our return.
Conclusion and Next Steps: Today’s EVA to Compass Rock was a blend of successful exploration and technological testing, tempered by the challenges of battery management and communication issues. These experiences provide valuable lessons for future missions, particularly in the areas of resource conservation and communication strategies. Encouraged by today’s accomplishments and equipped with new insights, we are optimistic about extending our exploratory capabilities in the upcoming EVAs.

EVA Report – November 30th

EVA #: 4

Today’s Date: 29 Nov 2023

Requested EVA Date: 30 Nov 2023

Requested start time for EVA: 12:00 PM

Requested end time for EVA: 03:00 PM

Does the weather report support EVA at this time?: Yes

Purpose of EVA: Reconnaissance. We want to also see if we can reach the Gateway to Lith with the rovers and on foot and come back within 3 hours. We understand there may be fossils around the area and duly acknowledge that they cannot be handled in any way.

Destination: Gateway to Lith

Coordinates (use UTM WSG 84): 518100;4255500

EVA Participants: Tungaa, Muggi, Dono

Road(s) and routes per MDRS Map: Marble Ritual -> Pooh’s Corner -> Cowboy Corner -> Gateway to Lith

Mode of travel: Driving and walking

Vehicles you will be using (If applicable): Curiosity and Perseverance

Supplemental Operations Report – December 2nd

Date: 12/2/2023
Name of person filing report: Ben Stanley
Reason for Report: Routine
Non-Nominal Systems: Director’s trailer heater, Perseverance parking brake not working

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 25%; Avg 76.7%.
VDC Last 24 hours: Max 59.34V; Min 40.44V; Avg 49.00V.
Generator run time: 4173.4 hours. Oil changed, bad part replaced, air intake manifold cleaned, locking washers added to air intake manifold bolts that were loose

Propane Readings:
Station Tank: 80%
Director Tank: 88%
Intern Tank: 88%
Generator Tank: 83%

Hab Static Tank – 264.3 gallons
GreenHab – 193.1 gallons
Outpost tank – 470 gallons
Science Dome – 0 gallons
Hab Toilet Tank emptied: Yes

Sojourner rover used: No.
Hours: 195.7
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: Perseverance parking brake non-functional

ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: Not used
General notes and comments: All cars awaiting renewed insurance ID.

Summary of Internet: All nominal.

EVA suits and radios:
Suits: All nominal
Comms: All nominal

Campus wide inspection, if action taken, what and why: Generator failed 11/28 and was repaired 11/30

Summary of Hab Operations: Nothing to report

Summary of GreenHab Operations: All nominal

Summary of SciDome Operations: All nominal

Summary of Observatories Operations: Robotic observatory opened during power failure. It’s closed now. Assisted Peter in resetting system 12/1

Summary of RAM Operations: All nominal

Summary of Outpost Operations: Director’s trailer heater is broken. New heater ordered and will be picked up in one week.

Summary of Health and Safety Issues: All nominal.

Operations Report – December 1st

SOL: 05

Name of person filing report: Munkh-Erdene "Muggi" Altankhuyag

Non-nominal systems: No non-operational system

Notes on non-nominal systems: n/a


Spirit rover used: the rover is used for EVA-05

Hours: 236 (before EVA)

Beginning charge: 100% (Before EVA)

Ending charge: 33% (On return from EVA, before recharging)

Currently charging: unplugged

Opportunity rover used: the rover is used for EVA-05

Hours: 146.4

Beginning charge: 100%

Ending charge: 35%

Currently charging: unplugged

Curiosity rover used: not used


Beginning charge:

Ending charge:

Currently charging:

Perseverance rover used: not used


Beginning charge: 100%

Ending charge: %

Currently charging: unplugged

General notes on rovers: The rovers have been taken to Somerville overlook.

Summary of Hab operations: Systems in the hab are operational.

WATER USE: 24.4 gallons

Water (static tank): 264.3 gallons

Static tank pipe heater (on or off): on

Static tank heater (On or off) on

Toilet tank emptied: yes

Summary of internet: Internet in the hab was disabled until 6:00 pm.

Summary of suits and radios: all radios are operational, and suits are operational but not charging, will be pulled in this night

Summary of GreenHab operations: Systems are operational.

WATER USE: 0.79 gallons

Heater: On

Supplemental light: will be operated from 10pm to 2 am

Harvest: 0.09lb

Summary of ScienceDome operations: The Dome systems are operational, EVA’s Cap Comm used the dome.

Dual split: Heat or AC, Off

Summary of RAM operations: RAM wasn’t used.

Summary of any observatory issues: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site